OpenStudy (anonymous):

Did I do this right?: √50=√5 x√10= 5√2

ganeshie8 (ganeshie8):

Looks good!

OpenStudy (anonymous):

@ganeshie8 for √32 is suppose to be √8 x√4 or is it suppose to be √4 x √8 ? Does it even matter?

ganeshie8 (ganeshie8):

it doesnt matter, but the easiest is to split it as \(\sqrt{16} \times \sqrt{2}\)

ganeshie8 (ganeshie8):

and recall that \(\sqrt{16}=4\)

ganeshie8 (ganeshie8):

so \[\large \sqrt{32}~=~\sqrt{16\times 2}=\sqrt{16}\times \sqrt{2} = 4\times \sqrt{2}=4\sqrt{2}\]

OpenStudy (anonymous):

Thank You ☺

ganeshie8 (ganeshie8):

np:)